Zeronknight (zeronknight):

Choose the inequality that is described below. Eight more than a number is greater than or equal to ten. A. x + 8 > 10 B. x - 8 > 10 C. x - 8 \geq 10 D. x + 8 \geq 10

Zeronknight (zeronknight):

C \[x - 8 \geq 10\]

Zeronknight (zeronknight):

D \[x + 8 \geq 10\]

OpenStudy (nethercreep333):

lol im too late

Zeronknight (zeronknight):

Don't give me a direct answer please, can you explain this?

OpenStudy (phi):

the point of this question is to see if you understand how to interpret English / equations for example Eight more than a number that should mean something to you, and the idea is to *automatically* understand it. to get the idea, think about this question what is 8 more than 2 ?

Zeronknight (zeronknight):

If it's what I think it is, you are adding 8 to 2.

OpenStudy (phi):

or even, what is 1 more than 2 ?

OpenStudy (phi):

in other words, in English 8 more than 2 means the same (as you guess) as 8+2 just as 1 more than 2 means 1+2 (i.e. 3)

OpenStudy (phi):

if they say 8 more than "a number" we give a name to "a number" (x is a favorite name) and using the same idea, we get 8 + x

OpenStudy (phi):

and because when you add two things, the order does not matter in other words 8+x is the same as x+8

Zeronknight (zeronknight):

Oh we want to figure out or isolate what X is?

OpenStudy (phi):

no , we don't really have to solve anything. We are translating Eight more than a number is greater than or equal to ten. into math so far we have 8+x is greater than or equal to ten ten (hopefully this is obvious) is written as 10 so 8+x is greater than or equal to 10 or (if this is multiple choice you have to know it might also be written this way) x+8 is greater than or equal to 10

Zeronknight (zeronknight):

How do we figure out if it is? And what do the underlines me on the >?

Zeronknight (zeronknight):

mean*

OpenStudy (phi):

the last thing to translate is "greater than" which is written > (which means the left side is bigger than the right side) or equal (put "half of an equal sign" under the > ) to get \( \ge\)

OpenStudy (phi):

the translation is \[ x+8 \ge 10\] we don't know if it is true or not. The only thing they are asking is how to change the words into that math.

OpenStudy (phi):

of course, once we have that "statement" we can do "math stuff" to it (i.e. simplify it) to \( x \ge 2 \) (that is the simplified version

Zeronknight (zeronknight):

oh! So the underline represents equal or more than at the same time, since X is unknown.. We have to put both = and > together?

OpenStudy (phi):

yes. there is a difference between being bigger than 10 versus being 10 or bigger

Zeronknight (zeronknight):

So the answer would be D, since we are adding, and we don't know if the problem is equal or greater than 10.

OpenStudy (phi):

also, it helps to remember that the "big" side of > goes next to the "big" side example 2 >1 is true (it says 2 is bigger than 1) and the big side of > is next to the bigger number

OpenStudy (phi):

yes. choice D is the answer. this is not a "math question" so much as a "can you rewrite English into math"

Zeronknight (zeronknight):

@phi Thank you for your help! Thank you for explaining to too so I can understand it! It's greatly appreciated! :)

OpenStudy (phi):

you should also learn what "1 less than 2" means in math it means 2-1 (and the answer is 1) 1 less than a number means x-1

Zeronknight (zeronknight):

Hmm. So I from what I understand. 2 less than 1 = 1-2 = 1-2= -1 and 1 less than 2 = 2-1=1

OpenStudy (phi):

exactly

Zeronknight (zeronknight):

Alright! Thanks for your help! :)

OpenStudy (phi):

for example 2 less than 1 in terms of temperature means 2 degrees colder than 1 degree means its -1 degree